OpenStudy (anonymous):

explain why the graphs of y=csc x and y=-ksinx will never intersect for any positive value of k

OpenStudy (anonymous):

@mml I think I can explain it, just give me a minute please.

OpenStudy (ranga):

y = csc(x) and y = -ksin(x) will intersect when csc(x) = -ksin(x) 1/sin(x) = - ksin(x) 1 = -ksin^2(x) sin^2(x) = -1/k The lefty hand side is always positive or zero. If k is positive, the right hand side will be negative and there will be no real solution.

OpenStudy (anonymous):

thanks

OpenStudy (ranga):

you are welcome.

OpenStudy (anonymous):

can you also explain what the Inverse Compotistion Rule is?

OpenStudy (ranga):

If f(x) is a function and g(x) is its inverse function then: f(g(x)) = x

OpenStudy (ranga):

The function f takes x and transforms it into f(x) The inverse function g takes f(x) and transforms it back to x. Or, the function g takes x and transforms it to g(x) The function f takes g(x) and transforms it back to x (that is, f(g(x)) = x).